### v2

# Projected sensitivity of CTAO to axion-like particles from blazars with a machine learning approach

Abstract
Blazars are a class of active galactic nuclei, supermassive black holes located at the centres of distant galaxies characterised by strong emission across the entire electromagnetic spectrum, from radio waves to gamma rays. Their relativistic jets, closely aligned to the line of sight from Earth, are a rich and complex environment, characterised by the presence of strong magnetic fields over parsec-scale lengths. Owing to their cosmological distance from Earth, these sources serve as ideal targets to probe non-standard gamma-ray propagation. In particular, axion-like particles (ALPs) could be detected through their coupling to photons, which enables ALP-photon conversions in external magnetic fields, leading to distinct signatures in the blazars’ gamma-ray spectra. In this work, we estimate the potential of the Cherenkov Telescope Array Observatory (CTAO) to constrain the ALP parameter space by simulating observations of two bright blazars, Mrk 501 and PKS 2155-304. We obtain projected $2\sigma$ exclusion regions, demonstrating that CTAO will be able to consistently improve present limits thanks to its greater energy resolution and point-source sensitivity with respect to present ground-based gamma-ray telescopes. In addition to the standard statistical technique based on the likelihood ratio test, we further demonstrate the application of a new method based on machine learning classifiers, which may help in reducing the effect of systematic model-dependent uncertainties in future ALP searches.
1 Introduction
Axion-like particles (ALPs) are a kind of hypothetical, pseudo-scalar elementary particles whose existence is predicted by several extensions of the Standard Model. Their name is derived from the original axion model, which was formulated by Peccei and Quinn Peccei and Quinn (1977), Weinberg Weinberg (1978), and Wilczek Wilczek (1978) as a solution of the strong CP problem in the context of quantum chromodynamics (QCD). While the existence of this original axion was soon ruled out Donnelly et al. (1978); Zehnder (1981), several “invisible axion” models were proposed, such as KSVZ Kim (1979); Shifman et al. (1980) and DFSZ Dine et al. (1981); Zhitnitsky (1980), which could elude experimental constraints while still solving the strong CP problem. In addition to these, it became clear that axion-like particles could be a more general prediction of beyond-the-Standard-Model theories such as Grand Unification Theories and string theories. These ALPs are not necessarily related to the original Peccei-Quinn model, but, like the QCD axion, they often arise as pseudo-Nambu-Goldstone bosons of spontaneously broken symmetry groups. Axions and ALPs are currently of great interest in cosmology and astrophysics, as they are among the best-motivated candidates as constituents of dark matter. Reviews of models and experimental searches for both QCD axions and ALPs can be found e.g. in refs. Jaeckel and Ringwald (2010); Ringwald (2012); Irastorza and Redondo (2018); Di Luzio et al. (2020).
The most important feature for the experimental detection of ALPs is their effective two-vertex coupling to photons, whose strength is controlled by the coupling constant $g_{a\gamma}$. Such a coupling is the basis for the most common approach to ALP searches, which exploits the expected ALP-photon mixing in the presence of external magnetic fields Sikivie (1983); Raffelt and Stodolsky (1988). This effect is employed, for example, in beam regeneration (or “light shining through walls”, LSW) experiments, such as ALPS-II Bähre et al. (2013), where a laser beam is shot against an opaque screen; in the presence of a strong enough magnetic field, some photons in the beam should be able to convert into ALPs, pass through the screen and oscillate back to photons on the other side. Other experiments for the direct detection of ALPs include helioscopes, like CAST Zioutas et al. (1999) and the foreseen IAXO Armengaud et al. (2019), and resonant cavity experiments such as ADMX Du et al. (2018) and MADMAX Brun et al. (2019).
As a general property of QCD axions, the axion-photon coupling $g_{a\gamma}$ is proportional to the axion mass $m_{a}$, as both quantities are inversely proportional to the Peccei-Quinn energy scale $f_{a}$ (e.g. Irastorza and Redondo (2018)). On the other hand, in the more general case of ALPs there is typically no fixed relationship between $m_{a}$ and $g_{a\gamma}$, so that they are independent coordinates spanning a 2-dimensional parameter space.
Astrophysical gamma-ray sources characterised by one or more magnetic field environments along their line of sight to Earth offer a natural framework to probe ALP-related effects, which has been extensively discussed in literature Hooper and Serpico (2007); Mirizzi et al. (2007, 2008); De Angelis et al. (2007, 2008, 2011); Horns et al. (2012); Batković et al. (2021); Galanti and Roncadelli (2022). Gamma-ray polarization effects could in principle observed, as ALPs only couple to the photon state parallel to the external transverse magnetic field; however, such effects are not detectable with present-day instruments Bassan et al. (2010); Galanti (2024). On the other hand, two main measurable spectral features may be investigated, whose extent generally depends on the field strength and topology, ALP parameters, source distance, and energy range of interest. The first is the presence of energy-dependent wiggles superimposed on the spectral shape, due to the peculiar shape of the photon survival probability $P_{\gamma\gamma}$ as a function of energy. The second is a spectral hardening at very high energy (VHE), due to gamma-ray photons converting into ALPs in the magnetic field of the source and back into photons in the Milky Way. This mechanism would allow the ALP-photon beam to travel cosmological distances, eluding the absorption that is normally experienced by VHE gamma rays due to the pair production (Breit-Wheeler scattering Breit and Wheeler (1934)) on the soft photons making up the extragalactic background light (EBL) Domínguez et al. (2011); Franceschini et al. (2008); Finke et al. (2010); Gilmore et al. (2012); Saldana-Lopez et al. (2021); De Angelis et al. (2013). The resulting effect, analogous to a “cosmic LSW experiment”, is expected to be more significant in the spectral tails of distant sources, for which the EBL absorption in the absence of ALPs would be strongest.
A typical target of these studies is represented by blazars. These sources are a particular class of active galactic nuclei (AGN), supermassive black holes characterised by accretion disks and strong relativistic jets at a small angle ($\lesssim 20\degree$) with the line of sight from Earth Padovani et al. (2017). Their broadband spectral energy distribution (SED) is characterised by a low-energy component due to the synchrotron radiation emitted by accelerated electrons in the jet, peaked between infrared and X-ray frequencies, and higher-energy emission in the gamma rays that may be produced either by leptonic or hadronic mechanisms Böttcher et al. (2013); Cerruti (2020). Leptonic mechanisms typically assume the gamma-ray emission to be originated by the same electron population that produces the synchrotron radiation via inverse Compton scattering, either on the synchrotron radiation field itself (synchrotron self-Compton, SSC) or on an external low-energy photon field (such as thermal photons from the broad line region, accretion disk or dusty torus, or the cosmic microwave background). In hadronic models, high-energy photons are emitted by a population of relativistic protons and other nuclei in the jet, typically via synchrotron radiation or photo-meson production, with the associated production of neutrinos. Based on their optical properties, blazars can be classified between BL Lacertae (BL Lac) objects and flat spectrum radio quasars (FSRQs). In addition, they can be distinguished between high-, intermediate- and low-synchrotron peaked (HSP, ISP and LSP) depending on their peak synchrotron frequency. Many blazars display significant spectral variability over timescales ranging from years down to minutes Böttcher (2019), with high-flux flaring states representing important opportunities of detection or even discovery of new sources. Despite being a very rare class of objects, almost all extragalactic sources detected in gamma-rays are associated with blazars.
The relativistic jets of blazars typically feature magnetic fields of order $\mathcal{O}(\rm mG)$ to $\mathcal{O}(\rm G)$ over $\mathcal{O}(\rm pc)$ scales. These objects may also be located inside galactic clusters with complex magnetic field structures of $\mathcal{O}(\rm\upmu G)$ intensity over $\mathcal{O}(\rm Mpc)$ scales; a notable example of this environment is represented by the Perseus galaxy cluster, which hosts the Seyfert galaxy NGC 1275. Additionally, being extragalactic sources, blazars are located at a cosmological distance from Earth (represented by the redshift parameter $z$), which makes EBL absorption effects important. All these characteristics may allow for hints of ALP-photon conversions in the gamma-ray spectra of blazars. Several constraints on the ALP parameter space have been obtained by analysing gamma-ray data from different blazars Abramowski et al. (2013); Jacobsen et al. (2023); Davies et al. (2023) and from the Perseus galaxy cluster Ajello et al. (2016); Abe et al. (2024a).
In this context, the next-generation Cherenkov Telescope Array Observatory (CTAO [^1] https://www.ctao.org/, formerly known as CTA) will be an ideal instrument to probe this kind of phenomena. The observatory will be made up of three different types of imaging atmospheric Cherenkov telescopes (IACTs), distributed in arrays over two sites in the Northern and Southern hemispheres Acharya et al. (2013). The Northern Array will be located at Observatorio del Roque de los Muchachos in La Palma, Canary Islands, and will be made up of 4 Large-Sized Telescopes (LSTs) and 9 Medium-Sized Telescopes (MSTs); the Southern Array, in Paranal, Chile, will feature 14 MSTs and 37 Small-Sized Telescopes (SSTs). These different configurations were designed to obtain the best sensitivity to sources visible from the two hemispheres, with the Southern array optimised to detect the highest energy photons (up to several hundreds of TeV) from the Galactic centre. In addition to this baseline “Alpha” configuration, an upgrade is foreseen by the Italian CTA+ Program [^2] https://pnrr.inaf.it/progetto-ctaplus/ to add 2 LSTs and 5 more SSTs to the Southern site Loporchio (2025). An alternative configuration has also been proposed, using telescopes of the Schwarzschild-Couder type Adams et al. (2021) to complement the standard MSTs (which feature a Davies-Cotton optical design). CTAO will obtain a significant improvement in all areas of performance with respect to present-days IACT systems: in particular, its greatly enhanced energy resolution and point-source sensitivity will make it an ideal instrument to study the spectra of distant sources in search of potential non-standard spectral features Acharya et al. (2018). The CTAO sensitivity over the ALP parameter space has been estimated by simulating observations of various blazars Meyer et al. (2014); Meyer and Conrad (2014) and of the Perseus galaxy cluster Abdalla et al. (2021).
Almost all of the previously cited works rely on the definition of some kind of test statistic (TS) based on the logarithm of the likelihood ratio (equivalently, the difference in log-likelihoods) to test the ALP hypothesis on real or simulated data Jacobsen et al. (2023); Davies et al. (2023); Ajello et al. (2016); Abe et al. (2024a); Meyer et al. (2014); Meyer and Conrad (2014); Abdalla et al. (2021). In mathematical terms, the TS is given by
${\rm TS}=-2\ln\frac{\mathcal{L}(H_{1})}{\mathcal{L}(H_{0})}\,$, (1)
where $H_{i}$ denotes a generic hypothesis and $\mathcal{L}$ is the likelihood function. The latter is defined as the probability that the model parameters take on certain values under $H_{i}$, conditioned to the observed data. In this work, we introduce a different approach for constraining the ALP parameter space with binned gamma-ray data, based on the use of machine learning (ML) classifiers. We apply our method to simulated CTAO observations of selected blazars, and compare the results with those obtained by using the standard likelihood-ratio test (LRT).
This paper is structured as follows. In Section 2 we describe the blazar selection, spectral modelling and simulation, as well as the modelling of ALP-photon oscillations. In Section 3 we review the standard LRT method to derive upper limits on the ALP parameter space, while in Section 4 we describe an alternative approach to the same problem based on the use of ML classifiers. In Section 5 we discuss and compare the results obtained with the two methods for the selected sources, and in Section 6 we present our conclusions.
2 Source modelling and simulation
2.1 Source selection
The observation of AGNs is one of the Key Science Projects (KSPs) proposed for CTAO, to be carried out within the first 10 years of operation of the observatory Acharya et al. (2018). The programme of this KSP will be organized in three parts: long-term monitoring of known AGNs, search and follow-up of flares, and measurement of high-quality spectra for different AGN classes and redshifts. One of the declared goals of the AGN KSP is to study possible imprints of new fundamental physics on the cosmological propagation of gamma-rays, including ALP searches Acharya et al. (2018); Abdalla et al. (2021).
In this work we shall focus on two of the most well-known and studied TeV blazars: Markarian 501 (Mrk 501) and PKS 2155-304. Both are high-synchrotron peaked BL Lac objects (HBLs), featuring significant emission at TeV-scale energies, and are included in the plans for long-term monitoring and high-quality spectral measurements of the KSP described above. This guarantees that CTAO will be able to gather a large amount of good quality data from these sources, which will be available for applying our proposed analysis.
The VHE spectral energy distribution (SED) of blazars is usually modelled with a simple function of energy, such as a power law (PL). Other common models include a log-parabola (LP)
[…]
which reduces to a simple power-law for $\beta\to 0$, or a power law with exponential cutoff (ECPL)
[…]
For the baseline states of our sources, we adopt the best-fit spectral models reported in the 4th Fermi-LAT Source Catalog Data Release 4 (4FGL-DR4) Abdollahi et al. (2022); Ballet et al. (2024). For Mrk 501 and PKS 2155-304, these are simple log-parabola functions. Since there are no observational constraints on the gamma-ray emission of blazars above a few tens of TeV, we conservatively assume that the validity of these models can be extended up to $10\,{\rm TeV}$ (thus not fully exploiting the CTAO energy range). In other works, an exponential cutoff factor around $1\,{\rm TeV}$ in rest-frame energy was added to these spectra (e.g. Abdalla et al. (2021)). We have checked that this modification does not significantly impact the extent of our final results.
As mentioned in the introduction, blazars are extragalactic objects, located at a distance measured by the cosmological redshift parameter $z$. For such targets, the absorption of gamma rays due to scattering on EBL photons must be taken into account. This effect is commonly quantified by an optical depth $\tau(E;z)$, such that the absorbed gamma-ray flux is attenuated by an exponential factor $\exp[-\tau(E;z)]$. Then, in the standard astrophysical scenario, the observed gamma-ray flux at Earth $\phi_{\rm obs}$ is related to the intrinsic gamma-ray flux at the source $\phi_{\rm int}$ by
$\phi_{\rm obs}(E)=\phi_{\rm int}(E)e^{-\tau(E;z)}\,.$ (4)
Since the effect of EBL absorption is negligible in the Fermi-LAT energy range, we take the 4FGL models as representative of the intrinsic spectra of the sources, and apply the EBL absorption factor to account for flux reduction at higher energies. Several models of the EBL exist in the literature, both phenomenological and obtained from first principles, mostly providing consistent results Domínguez et al. (2011); Franceschini et al. (2008); Finke et al. (2010); Gilmore et al. (2012); Saldana-Lopez et al. (2021); De Angelis et al. (2013). In the following, we shall employ tabulated values of the optical depth as computed in the model of Domínguez et al. Domínguez et al. (2011). We discuss the impact of choosing different EBL models on our final results in Appendix A.
In addition to the Fermi-LAT baseline states, we consider for each source a flaring state taken from previous VHE measurements. Specifically, we take the Mrk 501 flare observed by the MAGIC-I telescope on 30 June 2005 Albert et al. (2007), which is again well-fit by a LP shape, and the historical PKS 2155-304 flare observed by H.E.S.S. on 29 July 2006 Aharonian et al. (2009), which follows an ECPL. These spectra are corrected for EBL absorption and, as in the previous case, we assume their validity up to $10\,{\rm TeV}$ and we extrapolate them down to $30\,{\rm GeV}$.
The selection of spectral parameters for each source and activity state is reported and discussed in greater detail in Appendix B.
2.2 Modelling of ALP-photon oscillations
The propagating ALP-photon beam can be described as a three-state quantum system, representing the two photon polarizations and the ALP state, obeying a Schrödinger-like equation Raffelt and Stodolsky (1988). However, in the general case of an unpolarized beam (as for gamma rays), the system is more correctly represented by a $3\times 3$ density matrix $\rho$ satisfying the von Neumann-like equation De Angelis et al. (2011); Galanti and Roncadelli (2022)
[…]
where $y$ is the propagation direction and $\mathcal{M}_{0}(E)$ is an energy-dependent matrix whose elements account for photon dispersion, absorption and for ALP-photon mixing in external magnetic fields. For a given magnetic field configuration, this equation can be solved iteratively over discrete magnetic domains by the method of transfer matrices Mirizzi et al. (2008); Bassan et al. (2010); De Angelis et al. (2011); Galanti and Roncadelli (2022). Given the initial state $\rho_{i}$ at position $y_{0}$ and the total transfer matrix $\mathcal{U}(E;y,y_{0})$, the probability of finding the system in the final state $\rho_{f}$ at position $y$ is expressed as
[…]
In particular, for an initial unpolarized photon state $\rho_{0}=(1/2)\operatorname{diag}(1,1,0)$ and a final photon state of either polarization $\rho_{1}=\operatorname{diag}(1,0,0)$ or $\rho_{2}=\operatorname{diag}(0,1,0)$, the photon survival probablity is given by
[…]
In this work, the calculations outlined above are performed using the gammaALPs v0.3.0333https://github.com/me-manu/gammaALPs Python package Meyer et al. (2021). This package defines a ModuleList class where different magnetic field environments can be added in sequence. The ModuleList can then be run to automatically solve eq. (5) over the line of sight from a given source and obtain $P_{\gamma\gamma}(E)$ following eq. (7), for initial state $\rho_{0}$ and fixed ALP parameters $(m_{a},g_{a\gamma})$.
In order to compute the photon survival probabilities for our selected blazars, we consider ALP-photon mixing in their jets’ magnetic fields, modelled with helical and tangled components as in ref. Davies et al. (2021). The model is based on the blazar jet structure described by Potter and Cotter Potter and Cotter (2012a, b, 2013a, 2013b) and implemented in the gammaALPs class MixJetHelicalTangled. The parameters of each source are adapted from Figure 7 in ref. Potter and Cotter (2015) [^4] We note that ref. Potter and Cotter (2015) employs the same source nomenclature used in the the Fermi-LAT Bright Source List (0FGL) Abdo et al. (2009, 2010), and hence the source PKS 2155-304 appears listed as J2158., except for the free electron density normalization which is taken from Table 5 in ref. Tavecchio et al. (2010). Additionally, we include photon-ALP mixing in the Milky Way’s galactic magnetic field as modelled in ref. Jansson and Farrar (2012) and EBL absorption from ref. Domínguez et al. (2011). These propagation environments are implemented, respectively, in the classes MixGMF and OptDepth (from the ebltable package [^5] https://github.com/me-manu/ebltable). Finally, a possible contribution from the intergalactic magnetic field is usually modeled as a randomly oriented field with present-day intensity $B_{0}\lesssim 1\,{\rm nG}$ over a coherence length $l_{0}\lesssim 1\,{\rm Mpc}$ Ade et al. (2016). While the effect of this field is not necessarily small (e.g De Angelis et al. (2007); Abramowski et al. (2013)), we decided to neglect it owing to the large uncertainty in its structure and magnitude, in line with previous works Meyer and Conrad (2014); Ajello et al. (2016); Abdalla et al. (2021); Abe et al. (2024a).
In the presence of ALP-photon mixing, the intrinsic and observed fluxes are thus related by
$\phi_{\rm obs}(E)=\phi_{\rm int}(E)P_{\gamma\gamma}(E)\,$, (8)
where $P_{\gamma\gamma}$ reduces to the EBL absorption factor $e^{-\tau}$ for $g_{a\gamma}=0$. Figure 1 shows examples of the ALP effects on the baseline and flaring spectra of the blazars studied in this work.
Figure 1: ALP effects on the baseline and flaring spectral shapes chosen for the sources in this work. For large enough values of $g_{a\gamma}$, both spectral distortions (“wiggles”) and a hardening at high energies are visible
2.3 Data simulation and analysis
CTAO observations are simulated using the Gammapy v1.0.1 Python package Donath et al. (2023); Acero et al. (2023). For each source, we generate 50 hours of mock data for the baseline state and 5 hours (corresponding to about a night of observations) for the flaring state. We assume observations to be taken in the ON-OFF mode, i.e. by pointing the telescopes at an offset from the source (the ON region) and estimating the background from a number of OFF regions symmetrically distributed around the center of the field of view. For this setup, we consider an offset from the source of 0.7 degrees, taking ON and OFF regions with a radius of 0.2 degrees. We choose 3 OFF regions, or equivalently, an acceptance ratio $\alpha=1/3$. The source coordinates are implemented in the SkyCoord and SourceCatalog4FGL classes, respectively from the Astropy Astropy Collaboration and others (2013, 2018, 2022) and Gammapy [^6] Note that Gammapy v1.3 Acero et al. (2025) is necessary to load the 4FGL-DR4 catalog. packages. The parameters of flaring spectra are taken directly from refs. Albert et al. (2007); Aharonian et al. (2009), and the redshift parameters reported in TeVCat [^7] https://www.tevcat.org/ Wakely and Horan (2008) are used.
Table 1: Summary table of the astrophysical parameters of the selected blazars. The gammaALPs parameters are adapted from refs. Potter and Cotter (2015) and Tavecchio et al. (2010). Additionally, the parameters rvhe and r_T are used, both set equal to r_0 and taken to be $\sim 100$ times the size of the transition region $R_{T}$. Unspecified parameters are set to their default values
[…]
Table 2: Observation times, IRF file names and reference spectra for the selected blazars
Source Livetime [h] IRFs Reference spectrum Mrk 501 (baseline) 50 Prod5-North-20deg-AverageAz-4LSTs09MSTs.180000s LP Ballet et al. (2024) PKS 2155-304 (baseline) 50 Prod5-South-20deg-AverageAz-14MSTs37SSTs.180000s LP Ballet et al. (2024) Mrk 501 (flaring) 5 Prod5-North-20deg-AverageAz-4LSTs09MSTs.18000s LP Albert et al. (2007) PKS 2155-304 (flaring) 5 Prod5-South-20deg-AverageAz-14MSTs37SSTs.18000s ECPL Aharonian et al. (2009)
Figure 2: Example simulated SEDs for two 50-hour observations of the Mrk 501 baseline state. Spectral fits are shown for the ALP-less case and for two different ALP parameter sets. The residuals for both cases are computed as $(f-\phi)/\sigma_{f}$, where $f$ is the measured flux with uncertainty $\sigma_{f}$ and $\phi(E)$ is the model prediction (see ref. Abdalla et al. (2021))
We subdivide the measured energy axis with 20 bins per energy decade, in order to resolve the fine spectral irregularities that may be introduced by ALP effects. For each energy bin $i$, the expected number of signal counts $\mu_{i}$ is obtained by convolving the observed flux $\phi_{\rm obs}$ with the instrument response functions (IRFs), and then integrating over the energy bin width and observation time Abdalla et al. (2021). For our simulations we use the publicly available CTAO prod5 v0.1 IRFs Cherenkov Telescope Array Observatory and Cherenkov Telescope Array Consortium (2021), computed for the CTAO “Alpha” configuration described in Section 1. These include effective area, point spread function (PSF) and energy dispersion, as well as the background rates that can be used to compute the expected number of background counts per energy bin $b_{i}$. Different IRF files are available for 0.5, 5 and 50 hours of observations, both for the Northern and Southern Arrays in distinct subarray configurations, and for different values of zenith and azimuth angles. We simulate observations of Mrk 501 and PKS 2155-304 with the Northern and Southern Arrays, respectively, assuming all observations to be taken at a zenith angle of 20 degrees while averaging over the azimuth angle. The astrophysical and observational parameters used in simulating the datasets, as well as the chosen IRFs and spectral models, are summarised in Tables 1 and 2.
Datasets are simulated by sampling the signal and background counts in each energy bin from Poisson distributions, with average values $\mu_{i}+\alpha b_{i}$ in the ON region and $b_{i}$ in the OFF regions respectively, using the fake() method provided by the Dataset class in Gammapy. Example simulated SEDs for the baseline state of Mrk 501 are shown in Figure 2.
3 Derivation of ALP sensitivity with the likelihood-ratio test
Before presenting our method based on machine learning classifiers, we derive ALP limit projections using the standard LRT statistic, in order to introduce the general setup and to have a benchmark for our later results. Moreover, the results obtained with this approach can be directly compared with previous publications, where similar techniques were applied.
We consider an array of $10\times 10$ logarithmically spaced points over the ALP parameter space, with $m_{a}\in[0.1,1000]\,{\rm neV}$ and $g_{a\gamma}\in[0.03,7]\times 10^{-11}\,{\rm GeV}^{-1}$. The choice of the mass interval is commonly adopted in similar studies, and roughly corresponds to the region where the critical energy for the strong mixing regime falls within the CTAO energy range Meyer et al. (2014). Such energy value is defined as that above which the ALP-photon oscillation probability in a constant magnetic field becomes maximal and almost independent of energy Meyer et al. (2014); Mirizzi et al. (2008); Bassan et al. (2010); De Angelis et al. (2011); Galanti and Roncadelli (2022). On the other hand, the upper bound on $g_{a\gamma}$ is motivated by the most recent experimental limit set by the CAST helioscope Altenmüller et al. (2024).
Given the intrinsic spectrum of each source, we compute the photon survival probability $P_{\gamma\gamma}$ for each $(m_{a},g_{a\gamma})$ pair with gammaALPs as described in Section 2.2, and obtain the expected observed spectrum as in eq. (8). A likelihood function
$\mathcal{L}(m_{a},g_{a\gamma},\mu,b,\theta|D)=\prod_{i}\mathcal{L}(m_{a},g_{a\gamma},\mu_{i},b_{i},\theta|D_{i})$ (9)
is defined, where each $i$-th energy bin contributes with a Poissonian term of the form
$\begin{aligned} \mathcal{L}(m_{a},&g_{a\gamma},\mu_{i},b_{i},\theta|D)=\\ &\frac{(\mu_{i}+\alpha\,b_{i})^{n_{{\rm ON},i}}}{n_{{\rm ON},i}!}e^{-(\mu_{i}+\alpha\,b_{i})}\frac{b_{i}^{n_{{\rm OFF},i}}}{n_{{\rm OFF},i}!}e^{-b_{i}}\end{aligned}$ (10)
[…]
Following ref. Abe et al. (2024a), we then define the test statistic
${\rm TS}(m_{a},g_{a\gamma}|D)=-2\ln\frac{\mathcal{L}(m_{a},g_{a\gamma},\hat{\mu},\hat{b},\hat{\theta}|D)}{\hat{\mathcal{L}}(D)}\,$, (11)
where $\hat{\mathcal{L}}(D)$ is the maximum likelihood for the given dataset over the considered ALP parameter space. In practice, maximizing the likelihood is equivalent to minimizing the quantity
[…]
which corresponds to $2\ln\mathcal{L}$ after neglecting terms that do not depend on $\mu_{i}$ or $b_{i}$. This quantity is computed and returned by Gammapy when fitting each spectral model to the dataset $D$. Therefore, eq. (11) is equivalent to
$\begin{aligned} {\rm TS}(m_{a},&g_{a\gamma}|D)\\ &={\rm WSTAT}(m_{a},g_{a\gamma}|D)-{\rm WSTAT}_{\rm min}(D)\,.\end{aligned}$ (13)
Intuitively, this TS can be seen as a measure of the difference between any ALP model $(m_{a},g_{a\gamma})$ and the one that best fits the data.
Figure 3: Example likelihood-ratio TS distribution obtained for the baseline state of Mrk 501
For each source (considered in both baseline and flaring states) we simulate 100 observations without including ALP effects, following the prescriptions outlined in Section 2.3, and compute the average TS for each $(m_{a},g_{a\gamma})$ pair. In order to obtain the statistical significance for excluding the ALP hypothesis with parameters $(m_{a},g_{a\gamma})$, one then needs to compare the obtained average, denoted with ${\rm TS_{0}}(m_{a},g_{a\gamma})$, against the TS distribution expected under the ALP hypothesis. As has been argued several times, Wilks’ theorem Wilks (1938) cannot be applied to the case of ALPs Meyer and Conrad (2014); Abe et al. (2024a); Abdalla et al. (2021), and thus the shape of the TS distribution cannot be assumed a priori. Therefore, for each $(m_{a},g_{a\gamma})$ point and each considered source, we simulate 100 datasets including ALP effects and compute the corresponding TS values according to eq. (13). An example of a TS distribution obtained in this way is shown in Figure 3. Following ref. Abe et al. (2024a), we fit this distribution with a Gamma probability distribution function (PDF), given by
$f(x,\alpha,\beta)=\frac{\beta^{\alpha}x^{\alpha-1}e^{-\beta x}}{\Gamma(\alpha)}\,$, (14)
where $\alpha,\beta>0$ and $\Gamma(\alpha)$ is the gamma function
[…]
The confidence level (CL) for excluding the ALP hypothesis $(m_{a},g_{a\gamma})$ is then given by the cumulative distribution function (CDF) evaluated at ${\rm TS_{0}}(m_{a},g_{a\gamma})$, i.e.
[…]
Finally, the exclusion significance is expressed as the equivalent number of standard deviations ($z$-score)
$z=\sqrt{2}\,{\rm Erf^{-1}(CL)}\,$, (17)
where ${\rm Erf}^{-1}(x)$ is the inverse of the error function
[…]
All the statistical computations discussed above were performed using the Scipy Python package Virtanen et al. (2020).
Intuitively, it can be understood that for any $(m_{a},g_{a\gamma})$ pair the TS distribution will generally be concentrated around small values: since the best-fitting ALP model will most likely be close to the model that was used to generate the datasets, the difference between the two log-likelihood values (measured by the TS) should be small and only due to statistical fluctuations. Conversely, the average of ALP-less datasets may present much larger $\rm TS_{0}$} values for certain $(m_{a},g_{a\gamma})$ pairs, which in turn would allow them to be excluded. We stress here that, since we are testing datasets without ALP effects and we are computing the significance for excluding the ALP hypothesis, our results should be interpreted as an estimate of CTAO’s capability of constraining ALPs (i.e., placing upper limits on their parameter space) rather than discovering them.
4 Derivation of ALP sensitivity with machine learning classifiers
4.1 Machine learning classifiers
In this Section we give a brief review on the topic of ML classifiers, which we later apply in the derivation of upper limits on the ALP parameter space.
A basic example of a ML classifier is the binary decision tree (BDT). A BDT can act on a dataset [^8] In machine learning literature the term “sample” would be used, whereas “dataset” would denote a set of many samples., identified by a vector of features $\mathbf{X}=(X_{1},\dots,X_{n})$, and distinguish whether it belongs to one of two classes by operating sequential binary splits on the features’ values. The splitting thresholds at each step (node) are referred to as the parameters of the tree; on the other hand, quantities such as the tree depth, maximum number of splits, etc. are called hyperparameters. During the “supervised training” of the BDT, the splits are chosen in order to minimize some loss function, by operating on datasets whose class is known a priori. Once the tree has been trained and its parameters have been fixed, it can be used to classify previously “unseen” test datasets. There are several methods to evaluate the performance of a ML classifier during the training phase, e.g. evaluating its accuracy and related metrics with train-test splitting and cross-validation. This ensures that a similar performance will be achieved on new data as well. A vast literature of textbooks and papers exists on the many scientific applications of ML classifiers and BDTs; see e.g. ref. Raschka and Mirjalili (2017) for a basic introduction.
Random forests (RFs) are an example of ensemble classifiers, since, as the name suggests, they are based on averaging a large number of binary decision trees. In order to avoid overfitting, i.e. classifiers performing “too well” on training data but badly on test data, it is necessary that the single BDTs stop after few splits (it is said that they are “weak” classifiers). BDTs, RFs and several other ML classification algorithms are implemented in the open-source Python package Scikit-learn Pedregosa et al. (2011). When one of these algorithms is trained to distinguish between two classes and then applied to some sample, it typically returns a set of scores indicating how likely the given sample is to belong to each of the classes.
Machine learning classifiers such as RFs are widely used in Cherenkov telescope data analysis, specifically in steps that involve the so-called gamma-hadron separation Bock et al. (2004); Albert et al. (2008); Ohm et al. (2009). Here, the atmospheric shower images are parametrized by a set of quantities (the Hillas parameters Hillas (1985)) which are then used as features to distinguish between showers initiated by hadrons and by gamma-ray photons.
XGBoost (eXtreme Gradient Boosting) Chen and Guestrin (2016) is a similar algorithm to the RF, which leverages a technique called gradient boosting to efficiently minimise the loss function. Implementations of the XGBoost algorithm are available in several programming languages, including Python [^9] https://xgboost.readthedocs.io/en/stable/, and generally exhibit reasonably good performance “out of the box”, with minimal tweaking compared to RFs.
4.2 Method
In general, the effects of ALP-photon conversions on gamma-ray spectra are highly nonlinear and not easily parametrized. The energy and amplitude at which spectral features appear depend in a nontrivial way from the morphology and intensity of the magnetic fields encountered by the ALP-photon beam along the line of sight from Earth. It is therefore interesting to see how non-parametric statistical methods can address this task, such as those provided by ML classification algorithms. A first step in this direction is to verify that a ML-based method can reproduce the results obtained by the classical likelihood-ratio test (LRT) discussed above.
After being trained with a large amount of simulated spectra, both with and without ALP effects, a classifier is supposed to learn what the most important energy bins are to distinguish between the two classes, i.e., where the spectral oscillations and/or hardening are most relevant. This classification task crucially depends on the source under consideration, as well as on the ALP parameters $(m_{a},g_{a\gamma})$. Our approach was loosely inspired by that described in ref. Day and Krippendorf (2020).
For each of our selected sources and activity states, we define a grid of classifiers over the points of the ALP parameter space described in Section 3. The classifier corresponding to each $(m_{a},g_{a\gamma})$ pair is trained to distinguish between 20 000 simulated datasets including ALP effects and 20 000 only accounting for EBL absorption. The feature vector $\mathbf{X}$ used for training the classifiers is represented by the excess photon counts in each energy bin above background, which we assume to be already subtracted by conventional methods in the case of a real analysis.
These counts are then normalized to the $[0,1]$ interval, so that the only relevant characteristic for classifying the spectra is their shape, given by the intrinsic differences between adjacent bins rather than the overall number of counts. In order to allow for some variability in the intrinsic spectrum, avoiding to overfit the classifiers to a specific model, the training datasets are simulated by uniformly sampling values of each spectral parameter $p$ in the range $p-\Delta p,p+\Delta p]$, according to the values reported in Table 3. We have checked that the classifier performance only improves slightly by further increasing the number of training datasets.
The classifiers in the grid are based on the Python implementation of the XGBoost algorithm, using the following fiducial hyperparameters:
    n_estimators=100, max_depth=3,
    reg_lambda=500., reg_alpha=10.,
    eta=0.3
While this choice is arbitrary to a certain extent, we computed the values of several metrics (accuracy, precision, recall, F1 score) over the grid to check that the performance is reasonably good in the same regions of the parameter space where we expect ALP signatures to be stronger. Indeed, a decrease in performance is observed at large $m_{a}$ and small $g_{a\gamma}$ values, in the lower right corner of the grid, where the spectral distortions due to ALPs are weaker and less distinguishable from the ALP-less case. As an example, the results obtained for the baseline state of Mrk 501 are shown in Figure 4; see also Appendix C. These metrics were computed with Scikit-learn, using a 80-20 train-test split. For their definitions, see e.g. ref. Raschka and Mirjalili (2017).
Figure 4: XGB metrics for the baseline state of Mrk 501 over the ALP parameter space
Once the classifier grid for each source and spectral state has been trained, it is applied to 100 simulated datasets without ALP effects generated as described in Section 2.3. For each dataset $D$, the classifier corresponding to each $(m_{a},g_{a\gamma})$ pair will return a score $p_{\rm ALP}(m_{a},g_{a\gamma}|D)$ representing how likely the dataset is to be classified as ALP-like [^10] We note that in general this score is not a probability in a strict statistical sense, unless it is appropriately calibrated.. We then define the quantity
$\Pi(m_{a},g_{a\gamma}|D)=1-p_{\rm ALP}(m_{a},g_{a\gamma}|D)$ (19)
and, similarly to the previous case, we denote the average value obtained from the 100 ALP-less datasets with $\Pi_{0}(m_{a},g_{a\gamma})$.
We then search a $\Pi$ distribution for each $(m_{a},g_{a\gamma})$ pair, evaluating the $\Pi$ statistic in eq. (19) for 2 000 simulated ALP-like datasets. Since the value of $\Pi$ is always between 0 and 1, we fit the resulting distribution to a Beta PDF
$f(x,\alpha,\beta)=\frac{\Gamma(\alpha+\beta)x^{\alpha-1}(1-x)^{\beta-1}}{\Gamma(\alpha)\Gamma(\beta)}\,$, (20)
with $0\leq x\leq 1$ and $\alpha,\beta>0$. Using this distribution, the confidence level for excluding the ALP hypothesis and the $z$-score are computed using eqs. (16) and (17), respectively. The quality of the fit is generally good for medium and low values of $g_{a\gamma}$, as can be seen from Figure 5, but it worsens for higher ones. We have checked in Appendix D that this does not significantly impact the main conclusions of this work.
Figure 5: Example distributions of the $\Pi$ statistic defined in eq. (19), comparing good and suboptimal performance of a classifier algorithm in different regions of the ALP parameter space
Clearly, for a well-performing classifier, the $\Pi$ distribution obtained in this way should be largely concentrated around small values, i.e. an ALP-like dataset is likely to be classified correctly. On the other hand, the $\Pi_{0}$ value from the ALP-less datasets should be quite high (close to 1), thus leading to exclusion of the ALP hypothesis. There are, however, some regions of the ALP parameter space where the spectra with and without ALP effects are hardly distinguishable, typically for large $m_{a}$ and small $g_{a\gamma}$ values. The accuracy of the classifiers in these cases is then very low (cf. Figure 4), as can be seen from a $\Pi$ distribution peaked around 0.5, and no exclusion is expected. Example of distributions corresponding to these two situations are shown in Figures 5a and 5b, respectively.
Figure 6: ALP exclusion significance computed with the ML-based method, with the $2\sigma$ contours highlighted (solid lines). The contours obtained at the same confidence level with the LRT method are shown as dashed lines for comparison. A linear interpolation was applied between the points in the parameter space for which the exclusion confidence level was computed, which are indicated as black dots
5 Discussion of results
Our main results are presented in Figure 6, where a color map is used to represent the exclusion significance obtained with the ML classifier method in units of Gaussian standard deviations over the ALP parameter space. We note that, for both methods, the exclusion significance may be infinite for those points where the confidence level computed by eq. (16) is equal to 1. For plotting, these values are clipped to the highest finite significance; see also Appendix D.
The $2\sigma$ sensitivity obtained with the method of ML classifiers, shown as solid lines in Figure 6, is comparable with that resulting from the LRT (dashed lines). For all sources, the overall shapes and extensions of the projected exclusion regions are in agreement between the two methods, with occasional low-significance “holes” or high-significance “islands” in both sets of contours. Such features are common to other analyses of this kind Ajello et al. (2016); Abdalla et al. (2021) and do not significantly affect our main results.
The excluded region of the parameter space in the case of Mrk 501 is almost the same for both the baseline and flaring states. An explanation for this is the fact that the spectral shapes considered for the two states are fairly similar (see Fig. 1), and that the increased brightness of the flaring state is compensated by the longer exposure used for the baseline state. Given this, we can assume that the photon counts in the considered energy ranges are above the CTAO sensitivity for both states and are equally affected by ALPs, thus leading to similar exclusions in the parameter space. Conversely, PKS 2155-304 exhibits a spectral curvature and a stronger EBL absorption effect that significantly reduce the flux at higher energies. In this case, a flaring state can indeed bring the flux of higher-energy photons above the CTAO sensitivity, thus improving the constraining power of the observations. These qualitative considerations are confirmed by computing the Li-Ma significance $S$ of the signal Li and Ma (1983). While for Mrk 501 both observations yield $S\sim 300$, in the case of PKS 2155-304 a 5-hour flaring observation results in a significance $S\sim 800$, compared to $S\sim 200$ for a 50-hour observation in the baseline state.
These projections, representing the mean expected CTAO sensitivity to ALPs from each considered source, are obtained by averaging the results from many different simulated observations. A different approach for estimating such a mean sensitivity, employed in various studies on ALPs Meyer et al. (2014); Meyer and Conrad (2014); Abdalla et al. (2021) and WIMPs Acharyya et al. (2021, 2023); Abe et al. (2024b), would be to use so-called “Asimov datasets” Cowan et al. (2011) instead. These datasets are generated by simply setting the numbers of counts in the ON and OFF regions equal to their expected values, in order to eliminate Poissonian fluctuations and directly provide a median estimate of the sensitivity. We have checked explicitly that the contours obtained from a single Asimov dataset using the LRT method are very close to those resulting from averaging over many “noisy” datasets. On the other hand, we found that applying the ML classifiers to Asimov datasets causes a significant overestimation of the exclusion regions, as the total absence of Poissonian noise in photon counts almost always produces low $p_{\rm ALP}$ values even for extremely small $g_{a\gamma}$. In a realistic analysis, these points of the parameter space would not be excluded, as the inevitable photon noise could easily mimic ALP effects at very low couplings. We thus conclude that Asimov datasets cannot be reliably used in estimating ALP sensitivities with ML classifiers.
As a final remark, we note that these results may be affected by several systematic effects, such as the choice of energy binning, magnetic field parameters and EBL model. We discuss these effects in Appendix A. We note, in particular, that recent measurements of the cosmic infrared background by the CIBER experiment might extend the allowed parameter space for ALPs to a wider region than what is expected from the commonly adopted EBL models Kohri and Kodama (2017).
6 Conclusions
In this work, we have estimated the capability of CTAO to place upper limits on the ALP parameter space by simulating observations of the blazars Mrk 501 and PKS 2155-304 with the Northern and Southern Arrays of the Observatory. The source selection and observation criteria were motivated by the AGN Key Science Project described in ref. Acharya et al. (2018). Mock datasets for each source were prepared considering both long-term observations of baseline states, with spectra taken from the Fermi-LAT 4FGL-DR4 source catalog Ballet et al. (2024), and short flares based on previous TeV observations Albert et al. (2007); Aharonian et al. (2009).
In Figure 7 we compare the projections derived in this work with previously published constraints on the ALP parameter space [^11] The reference limits are mostly taken from the online repository at ref. O’Hare (2020).. Most of these are computed at the $2\sigma$ confidence level Dessert et al. (2022); Noordhuis et al. (2023); Manzari et al. (2024); Ajello et al. (2016); Altenmüller et al. (2024); Abdalla et al. (2021); Abramowski et al. (2013), except for the limits by MAGIC which are reported at the $3\sigma$ level Abe et al. (2024a). In particular, ref. Abdalla et al. (2021) represents the only previous estimate of ALP sensitivity officially published by the CTAO Consortium, derived from simulated observations of the AGN NGC 1275 inside the Perseus galaxy cluster. With these targets, we expect that CTAO will be able to place limits on a much wider portion of previously unconstrained ALP parameter space in the $0.1\mbox{--}100\,\rm{neV}$ mass range. We finally remark that, while already competitive, the limits derived from a Southern source like PKS 2155-304 could be further improved with the foreseen addition of 2 LSTs to the Southern Array of CTAO, whose contribution would increase the sensitivity to low-energy spectral oscillations.
Figure 7: Comparison of the CTAO $2\sigma$ exclusion regions on the ALP parameter space obtained in this work with reference constraints from the literature (green). The projected sensitivities obtained from 50-hour observations of Mrk 501 (red) and PKS 2155-304 (blue) in their baseline states, as well as that from a 5-hour observation of a PKS 2155-304 flare (yellow) are shown. We do not explicitly include the results obtained from the Mrk 501 flare as they largely overlap with those from the baseline observation
Astrophysical ALP searches are affected by large systematic errors. These mainly stem from the uncertainty in the modeling of the morphology and intensity of the magnetic fields along the line of sight from the source. The magnetic field structure in blazar jets, in particular, is almost unconstrained experimentally and can only be estimated from theoretical arguments, thus making ALP searches from blazars somewhat model-dependent. In the absence of a more precise determination of this structure, turning to different computational approaches that could detect ALP-induced spectral features in a model-independent fashion might be a viable strategy. To probe this possibility, we computed CTAO sensitivity projections by applying the classical likelihood-ratio test and a newly proposed method making use of machine-learning classifiers. While in the present context the classifier-based method uses the same modeling assumptions as the traditional method, thus not really solving the issue of systematic uncertainties, a promising result is shown by the fact that the sensitivities obtained by the two results broadly agree with each other, as shown in Figure 6. This is a first step in the direction of exploring different non-parametric methods for ALP searches in astrophysical data. Future developments of this concept will involve dedicated studies focused on the application of more complex and optimized algorithms, such as neural networks, with the aim of maximizing the attainable experimental sensitivity to ALPs while reducing the amount of model-dependent assumptions.
We conclude that observations of blazars with CTAO will be a powerful means to probe a significant portion of the parameter ALP space beyond the regions previously excluded by experiment, helping in guiding searches for these elusive particles as part of the AGN Key Science Project. To this aim, in this work we have introduced an alternative approach to constrain the ALP parameter space from binned gamma-ray data, based on the use of machine-learning classifiers, which offers promising results in good agreement with the standard likelihood-ratio test.
Acknowledgments
We thank the CTAO internal reviewers, G. D’Amico, G. Galanti and M. Garczarczyk, for their insightful suggestions. F. S. would additionally like to thank A. Di Florio, F. Simone, F. Cuna, A. Mirizzi and K. Kohri for the comments they provided during various phases of this work.
This work was conducted in the context of the CTAO DMEP working group. CTAO gratefully acknowledges financial support from the agencies and organizations listed at https://www.ctao.org/for-scientists/library/
acknowledgments/
. F. S. acknowledges financial support by the EU funding program “NextGenerationEU” in the context of the PNRR-IR “CTA+ - Cherenkov Telescope Array Plus” program.
We would like to thank the computing centers that provided resources for the IRF generation, listed at
https://zenodo.org/records/5499840.
Appendix A Systematic effects
In this Appendix we discuss how our results are affected by changes in energy binning, magnetic field strength or EBL model. While the former effects are similar between the sources considered in this work, we focus our discussion on PKS 2155-304 as it is the most strongly affected by EBL absorption.
Figure 8a shows the average $2\sigma$ sensitivity on the ALP parameter space from a 50-hour observation of PKS 2155-304 in its baseline state for different choices of the energy binning. We compare the results presented in the main text, obtained for 20 bins per decade, with those resulting from considering either half or double the number of bins per decade (10 or 40, respectively). The latter choice of binning, in particular, was employed in ref. Abdalla et al. (2021). While there is very little difference between these results, halving the number of bins seems to reduce the size of the hole-like features around $g_{a\gamma}\sim 4\times 10^{-11}\,{\rm GeV}^{-1}$, while doubling it turns the feature into a continuous band. A similar outcome is found by comparing the results from different EBL models Franceschini et al. (2008); Finke et al. (2010); Gilmore et al. (2012) to the one used in the main text Domínguez et al. (2011), with the sensitivity to ALP parameters almost unaffected by this choice (Figure 8b).
We consider in Figure 8c how our results are affected by changing the average magnetic field strength $B_{0}$ in the jet, either increasing it by 50% ($B_{0}={1.23\,\rm G}$) or decreasing it by the same amount ($B_{0}={0.41\,\rm G}$) with respect to the value reported in Table 1 ($B_{0}={0.82\,\rm G}$). As the ALP effects on the photon spectrum depend on the product between the coupling constant $g_{a\gamma}$ and the magnetic field, it is easy to see that increasing the latter allows to set more stringent limits on the former, and vice versa. We remark that this uncertainty comes not only from the large experimental difficulty in determining the value this parameter, but also from the fact that such value is unlikely to remain constant while the source goes through different activity states.
Figure 8: CTAO sensitivity to the ALP parameter space obtained from a 50-hour simulated observation of PKS 2155-304 in its baseline state for different choices of the energy binning, EBL model, magnetic field strength and training strategy
Finally, in Figure 8d we compare our main results with those we find by not randomizing the intrinsic spectral parameters within their error bands in the training datasets (see Section 4.2) and only using their average values. As expected, the sensitivity derived in this way is stronger, as the ML classifiers become overfit to a specific spectral model, but less generalizable. On the other hand, in a “randomized” training, the variations on the spectral parameters should not be larger than a few percent in order to maintain a solid performance of the classifiers.
Appendix B Details on spectral models
The parameters of each spectral model considered in this work are reported in Table 3, where LP and ECPL refer to the log-parabola and power law with exponential cutoff models defined in eqs. (2) and (3). These models were implemented in Gammapy using the LogParabolaSpectralModel and ExpCutoffPowerLawSpectralModel classes, respectively. As discussed in the main text, the baseline spectra were taken from the Fermi-LAT 4FGL-DR4 catalog Ballet et al. (2024), while the flaring spectra were selected from previously published IACT observations. In particular, the EBL-corrected parameters of the Mrk 501 flaring spectrum are taken from Table 7 and Figure 19 of ref. Albert et al. (2007), referring to the flare of 30 June 2005. The PKS 2155-304 parameters are quoted from Table 3 in ref. Aharonian et al. (2009) and correspond to the data taken with the lowest energy threshold ($\rm 200\,GeV$, labelled as T200), corrected for EBL absorption.
Table 3: Parameters of the spectral models used for the sources considered in this work. We note that the $\beta$ parameter for the Mrk 501 flaring state is scaled by a factor $\ln(10)$ with respect to ref. Albert et al. (2007) after the conversion from base-10 logarithm to natural logarithm
Source Model $\phi_{0}$ $E_{0}$ $\alpha$ $\beta$ $E_{\rm cut}$ Ref. Baseline $10^{-12}\,\rm{MeV^{-1}\,cm^{-2}\,s^{-1}}$] [MeV] Mrk 501 LP $3.76\pm 0.05$ 1535.09 $1.75\pm 0.01$ $0.019\pm 0.004$ - Ballet et al. (2024) PKS 2155-304 LP $2.59\pm 0.13 11$60.97 $1.77\pm 0.01$ $0.042\pm 0.004$ - Ballet et al. (2024) Flaring $10^{-10}\,\rm{TeV^{-1}\,cm^{-2}\,s^{-1}}$] [TeV] [TeV] Mrk 501 LP $18.6\pm 0.06$ 0.3 $1.73\pm 0.06$ $0.13\pm 0.03$ - Albert et al. (2007) PKS 2155-304 ECPL $4.51\pm 0.36$ 1.0 $2.16\pm 0.05$ - $1.74\pm 0.24$ Aharonian et al. (2009)
Figure 9: Accuracy of the XGBoost classifier grids defined over the ALP parameter space with the hyperparameters defined in Section 4, with the $90\%$ contours highlighted
Appendix C Classifier accuracy over the ALP parameter space
In Figure 9 we report the accuracy of the XGBoost classifiers defined over the ALP parameter space for each source, using the hyperparameters defined in Section 4. As noted in the main text, we observe that the classifiers’ performance is reasonably good (accuracy $\gtrsim 90\%$) in the same regions of the parameter space where we expect CTAO to cast the strongest limits. All the other metrics that we have checked (precision, recall, F1 score) have almost the same contours as accuracy, similarly to Figure 4.
Appendix D Goodness of fit for the Beta distribution
The example distributions of the $\Pi$ statistic shown in Figure 5 refer to cases where the accuracy of the classifiers is either worst or average. In these situations, the $\Pi$ values are more or less spread across the [0,1] interval and the resulting distribution is typically well fit by the Beta distribution of eq. (20). This can be checked e.g. by applying the Kolmogorov-Smirnov (KS) test [^12] https://docs.scipy.org/doc/scipy/reference/generated/
scipy.stats.kstest.html
, which returns a large $p$-value when the data are compatible with the proposed distribution (i.e., the hypothesis that the data are described by said distribution is not rejected).
In cases where the classifier accuracy is maximal, i.e. for large $g_{a\gamma}$ values, the resulting $\Pi$ distribution is largely peaked around very small values, as discussed in Section 4 (see e.g. Figure 10). In such extreme occurrences, the fit to the Beta distribution is typically not as good (resulting in small $p$-values from the KS test), along with being very sensitive to the numerical precision of the data. Moreover, the average value $\Pi_{0}$ obtained from the ALP-less datasets is usually close to 1, resulting in a confidence level for excluding the ALP hypothesis of essentially 100%, and consequently in a divergent $z$-score; see eqs. (16) and (17).
In order to determine whether our main results hold under these limitations, we recomputed the exclusion confidence levels for each point in the ALP parameter space simply as the fraction of the 2000 simulated $\Pi(m_{a},g_{a\gamma})$ values smaller than $\Pi_{0}(m_{a},g_{a\gamma})$. This approach is in fact equivalent to eq. (16), but it does not require to assume a specific distribution for $\Pi$. By doing so, we have explicitly checked that the 95% CL ($2\sigma$) sensitivity region on the ALP parameter space remain unchanged. On the other hand, without the smoothing provided by a fitting function, the $z$-score diverges to infinity much more quickly as $g_{a\gamma}$ increases. We conclude that higher-confidence limits (e.g. $3\sigma$, $5\sigma$) obtained with the classifier-based method should be regarded as less reliable and more dependent on the quality of the fit.
Figure 10: Example of a highly skewed $\Pi$ distribution obtained for a ML classifier close to maximal accuracy